Silver Company makes a product that is very popular as a Mother’s Day gift. Thus, peak sales occur in May of each year, as shown in the company’s sales budget for the second quarter given below:

April May June Total
  Budgeted sales (all on account) $300,000   $500,000   $200,000   $1,000,000

     From past experience, the company has learned that 20% of a month’s sales are collected in the month of sale, another 70% are collected in the month following sale, and the remaining 10% are collected in the second month following sale. Bad debts are negligible and can be ignored. February sales totaled $230,000, and March sales totaled $260,000.

Required:
1. Prepare a schedule of expected cash collections from sales, by month and in total, for the second quarter.

2. Assume that the company will prepare a budgeted balance sheet as of June 30. Compute the accounts receivable as of that date.
Answer #1

Schedule of expected cash collection

April May June Total
February Sales 230000*10% = 23000 23000
March Sales 260000*70% = 182000 26000 208000
April Sales 60000 210000 30000 300000
May Sales 100000 350000 450000
June Sales 40000 40000
Total Cash collection 265000 336000 420000 1021000

Account receivable = (500000*10%+200000*80%) = 210000